OpenStudy (anonymous):

(-8y^8/5x^4)^-3 simplify

OpenStudy (bibby):

properties of exponents \((\dfrac{a}{b})^M=\dfrac{a^M}{b^M}\\ \)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

u wot m8

OpenStudy (bibby):

u w0t?

OpenStudy (anonymous):

wot

OpenStudy (bibby):

\(\large (\dfrac{-8y^8}{5x^4})^{-3}=\dfrac{(-8y^8)^{-3}}{(5x^4)^{-3}})\)

OpenStudy (anonymous):

oh sorry ue positive exponents to simplify

OpenStudy (anonymous):

use

OpenStudy (bibby):

well we can simplify the bottom because \(\dfrac{1}{a}=a^{-1}\)  sorry internet disconnected